# Measuring Dwarf Galaxy Intrinsic Abundance Scatter with Mid-resolution Spectroscopic Surveys: Calibrating APOGEE Abundance Errors

### v1

@@ Abstract
The first generations of stars left their chemical fingerprints on metal-poor stars in the Milky Way and its surrounding dwarf galaxies. While instantaneous and homogeneous enrichment implies that groups of co-natal stars should have the same element abundances, small amplitudes of abundance scatter are seen at fixed [Fe/H]. Measurements of intrinsic abundance scatter have been made with small, high-resolution spectroscopic datasets where measurement uncertainty is small compared to this scatter. In this work, we present a method to use mid-resolution survey data, which has larger errors, to make this measurement. Using APOGEE DR17, we calculate the intrinsic scatter of Al, O, Mg, Si, Ti, Ni, and Mn relative to Fe for 333 metal-poor stars across 6 classical dwarf galaxies around the Milky Way, and 1604 stars across 19 globular clusters. We first calibrate the reported abundance errors in bins of signal-to-noise and [Fe/H] using a high-fidelity halo dataset. We then apply these calibrated errors to the APOGEE data, and find small amplitudes of average intrinsic abundance scatter in dwarf galaxies ranging from 0.032 -- 0.14 dex with a median value of 0.043 dex. For the globular clusters, we find intrinsic scatters ranging from 0.018 -- 0.21 dex, with particularly high scatter for Al and O. Our measurements of intrinsic abundance scatter place important upper limits on the intrinsic scatter in these systems, as well as constraints on their underlying star formation history and mixing, that we can look to simulations to interpret.
@@ 1 Introduction
Quenched dwarf galaxies that are bound to and inhabit our stellar halo serve as laboratories for studying snapshots of the early chemical evolution of the universe. With average stellar metallicities typically $\rm{[Fe/H]}<-1$ (Kirby et al., 2008, 2013; Geha et al., 2013; Li et al., 2018; Fu et al., 2023), their oldest, most metal-poor stars contain the chemical fingerprints of the very first generations of stars, including Population III (Frebel & Bromm, 2012; Ji et al., 2015; Hartwig et al., 2018, 2019). An understanding of the contribution to chemical enrichment from Population II stars is also necessary for placing constraints on the chemical evolution of the early universe. In all, to decode the signatures of the chemical ancestry of dwarf galaxies, it is essential to understand the observational effects of the propagation of elements from various populations of stars and from a variety of nucleosynthetic sources throughout the ISM (Ji et al., 2015).
@@
The chemical history of a stellar population is observed through measurements of element abundances, as well as their intrinsic abundance scatter – the true scatter of the data, accounting for measurement uncertainties – which necessitates careful consideration of uncertainties on abundances. Previous studies have uncovered overall element abundance scatter in numerous abundances, both within and between dwarf galaxies (Hill et al., 2019; Ji et al., 2020, 2022), but there remains a need to resolve the intrinsic scatters of chemical abundances above the level of observational errors; a difficult task for even small high-resolution spectroscopic studies. Mid-resolution spectroscopic surveys [e.g. APOGEE (Majewski et al., 2017), GALAH (De Silva et al., 2015), Gaia RVS (Gaia Collaboration et al., 2016, 2023) and LAMOST (Cui et al., 2012a, b)] afford the community with copious amounts of data, and target numerous stellar systems. Nonetheless, measurements of the intrinsic abundance scatter are made even more challenging by larger measurement uncertainties than higher signal-to-noise (SNR), higher resolution data, as well as potentially inaccurate estimates of measurement uncertainty in some parameter spaces. An important question to ask, however, is if we can use large mid-resolution surveys, to test the the intrinsic scatter of metal-poor systems that have been targeted and furthermore if we can validate or if necessary, calibrate, the uncertainty estimates so that these are accurate.
@@
From a theoretical perspective, the intrinsic scatter of element abundances in any system of stars is dependent on a number of properties. This includes the efficiency of gas mixing in the ISM (Krumholz & Ting, 2018; Emerick et al., 2020a; Ji et al., 2023), metal-retention and recycling through inflows and outflows (Emerick et al., 2018, Mead et al. in prep), stochastic star formation (Welsh et al., 2021; Pan & Kravtsov, 2023; Griffith et al., 2023), and mass and metallicity dependence of nucleosynthetic sources (Muley et al., 2021). The assessment of the scatter of individual elements is one indirect approach to access these environmental parameters.
@@
Because different nucleosynthetic channels contribute to the enrichment of different elements, intrinsic scatter is also a metric by which we can study the contribution of different nucleosynthetic sources to the chemical evolution of different stellar systems, as well as the mass and metallicity dependence of yields. For example, light elements such as C, N, and F are primarily produced by low and intermediate mass stars on the asymptotic giant branch (AGB) (Karakas, 2010; Kobayashi et al., 2011). The $\alpha$-elements (e.g. O [^1] O is also produced through the CNO cycle, and so not always considered an $\alpha$-element, but this contribution is subdominant for low-metallicity stars., Mg, Si, Ca, Ti) are primarily produced by massive stars and ejected by core-collapse (Type II) supernovae (SNe) (Timmes et al., 1995; Kobayashi et al., 2006). About half of Fe-peak elements (e.g. Cr, Mn, Fe, Ni, Co) are produced in Type Ia supernovae (Kobayashi & Nomoto, 2009). The well-known [$\alpha$/Fe]-[Fe/H] relationship, which shows the delay in enrichment in Fe from Type Ia SNe versus the short-timescale production of $\alpha$-elements by Type II SNe, can be used to characterize the star formation rates in galaxies (e.g. Tinsley, 1980; Matteucci & Greggio, 1986; Hayden et al., 2015).
@@
Other elements, such as the odd-Z elements (e.g. Al, P) and neutron-capture elements (s- and r-process elements) depend on the excess of neutrons available. Odd-Z elements depend on the metallicity of the progenitor star as their formation requires a surplus of neutrons made from ${}^{22}$e during He-burning (Kobayashi et al., 2011). Strong s-process elements (Sr to Pb) are produced in low-mass AGB stars (Busso et al., 1999; Herwig, 2005; Karakas & Lattanzio, 2014), whereas weak s-process elements (between Fe and Sr) are produced in massive stars near solar metallicity (Pignatari et al., 2010). The formation mechanism for r-process elements is debated but neutron star mergers are a likely scenario (Lattimer & Schramm, 1974; Rosswog et al., 1999; Goriely et al., 2011; Wanajo et al., 2014).
@@
Kobayashi et al. (2020) developed galactic chemical evolution models to place constraints on the origin of elements as a function of time and galactic environment. Models such as this, along with an understanding of gas mixing, the ratio of nucleosynthetic sources, and stochasticity of star formation can be used to predict the origin and amount of scatter in dwarf galaxy observations. High-resolution star-by-star models with individual stellar feedback (e.g. Brauer et al. in prep; Andersson et al. in prep) are uniquely poised to explore these physical drivers in the context of intrinsic abundance scatter.
@@
In the solar neighborhood, Griffith et al. (2023, hereafter G23) measured element abundances and intrinsic abundance scatters for metal-poor stars for 12 elements ranging from 0.04 for [Cr/Fe] to 0.16 for [Na/Fe] dex. Because of their low metallicity, these stars are likely field halo stars. They notably show that these scatters can be produced by Nsimilar-$\sim$50 core-collapse supernovae, which is a number too small to accurately represent an IMF.
@@ 2.1 Data Cuts
For all objects, we implement a metallicity cut of $-2.1\leq$ [Fe/H] $\leq-1$ to obtain the metal-poor stars whose scatter we are interested in. For all but the halo stars, we identify likely members of each system using the assigned APOGEE MEMBERFLAG. Halo stars are identified as metal-poor stars with distance $d\leq 2500\rm{pc}$. As in G23, these stars are from the solar neighborhood, but due to their low-metallicity, are likely halo stars. We remove stars in known substructure including Gaia-Sausage-Enceladus (GSE) (Horta et al., 2023a) and globular clusters (Schiavon et al., 2024) from this sample. We also tested a stricter cut on GSE stars by eliminating high eccentricity (e $>$ 0.7) stars, but this had no effect on our results.
To ensure the quality of the data, we make additional data cuts. We remove all stars which have been flagged with warnings generated by ASPCAP for their stellar parameters (STAR_BAD) or abundance measurements (M_H_BAD; ELEMFLAG). To exclude very low-quality spectra, we only use stars with a $\rm{S/N}\geq 25$ for our analysis, with the final data cut for dSphs having a median $\rm{S/N}=61.22$. Lastly, we only include stars with $4000\leq\rm{T_{eff}}\leq 5000K$ (median $\rm{T_{eff}}=4434\,\rm{K}$ for dSphs) and $0.0\leq\rm{log}g\leq 3.0$ (median $\rm{log}g=0.99$ for dSphs), keeping the range wide enough to contain as much of the data as possible but narrow enough to avoid systematic variations of these quantities with [X/Fe] (e.g. Griffith et al., 2021; Weinberg et al., 2022). Table 3 in the Appendix includes a summary of the resulting number of stars in our sample for each object we study. Because Bootes I has so few metal-poor stars at the required SNR, we exclude Bootes I from analysis.
@@ 2.2 Data Summary
We use chemical abundances derived by ASPCAP for the stars in our sample (García Pérez et al., 2016). We focus on the following elements: Al (light elements); O, Mg, Si, Ti ($\alpha$-elements); and Ni and Mn (Fe-peak elements). Figure 1 shows the abundance [X/Fe] and metallicity [Fe/H] measurements for stars in our dSph sample. Data points in Figure 1 are colored by $\rm{T_{eff}}$ to demonstrate that there are no trends in [X/Fe] with $\rm{T_{eff}}$ and that measured scatter is independent of stellar parameters. Generally, we see negative trends in [X/Fe] with [Fe/H]. As we are interested in processes that influence scatter in abundances for stars at the same metallicity, or in other words, the scatter around the trend line, we fit and subtract a second order model to the data in the [Fe/H]-[X/Fe] plane. In Figure 2, we show the resulting residuals for the dSphs, revealing the scatter of our sample with [Fe/H]. We similarly do this for the Milky Way stellar halo and GC stars in our sample. This trend-subtracted data is used in all further analysis.
Figure 3: Example line windows for star 2M02383346-3428246 in Fornax. The combined spectrum is shown in black with the ASPCAP model in red. The line window is indicated by the green shading.
@@ 2.3 Spectra Quality Check
Metal-poor stars have weak metal lines which make abundance measurements uncertain. Therefore, we directly examine the spectra of stars in our sample to assess the quality of the ASPCAP fit for spectral lines of elements we are using. We use the strongest spectral lines from the line list (Smith et al., 2021) used in the APOGEE abundance pipeline. Ultimately, we find that Al, O, Mg, Si, Ti [^2] Ti is labeled as ‘deviant’ from literature expectations in DR17, however visual inspection shows the lines are well-fit for the stars in our sample., and Ni show absorption features, which are generally well-fit by the model. We conclude that it is reasonable to proceed with our analysis using these abundance measurements. We also include Mn in our analysis, but note that these lines were not as well-fit, and that these abundance measurements should be treated with caution and assumed to be upper limits. Figure 3 shows a typical example of ASPCAP line fits for the stars and elements included in our sample. In addition to a visual inspection, we implement a quality cut whereby we calculate the $\chi^{2}$ for the ASPCAP best fit model (i.e. $\chi^{2}=\Sigma\frac{(\rm{x-y})^{2}}{\rm{\delta x}^{2}}$) and remove individual abundance measurements with a reduced $\chi^{2}>6$ from consideration in the analysis.
@@ 3.1 APOGEE Data Intrinsic Scatter
APOGEE has errors on the level of intrinsic scatter reported in other literature, therefore we compare the average scatters calculated with the APOGEE data used in our work to average scatters of other data for dwarfs in the literature, namely, the Kirby et al. (2018, hereafter K18) dataset which carefully accounts for the random and systematic uncertainties. From this data set, we only consider dSphs with at least 10 observed stars. This dataset overlaps partially with the APOGEE dataset (Draco, Fornax, Sculptor, Sextans, and Ursa Minor) and includes 3 additional dSphs (Canes Venatici I, Leo I, and Leo II). Figure 4 also shows that the uncalibrated data for the dSphs in the APOGEE dataset have higher average intrinsic scatter than the average for those in K18. This difference is substantial, and ranges from 0.02 dex to 0.1 dex.
Similarly high intrinsic abundance scatters can be shown for field stars in the APOGEE halo data when compared to intrinsic abundance scatters calculated in the same way from G23 data. In Figure 5, we report the intrinsic abundance scatters in the halo for the same elements. These range from a minimum of 0.09 dex for [Ni/Fe] to a maximum of 0.20 dex for [Al/Fe]). We also show that these values range from 0.02 up to 0.09 dex higher than intrinsic scatters for halo stars in G23. These differences call into question the accuracy of the APOGEE uncertainty estimates for the metal-poor stars.

### v2

@@ Abstract
The first generations of stars left their chemical fingerprints on metal-poor stars in the Milky Way and its surrounding dwarf galaxies. While instantaneous and homogeneous enrichment implies that groups of co-natal stars should have the same element abundances, small amplitudes of abundance scatter are seen at fixed [Fe/H]. Measurements of intrinsic abundance scatter have been made with small, high-resolution spectroscopic datasets where measurement uncertainty is small compared to this scatter. In this work, we present a method to use mid-resolution survey data, which has larger errors, to make this measurement. Using APOGEE DR17, we calculate the intrinsic scatter of Al, O, Mg, Si, Ti, Ni, and Mn relative to Fe for 333 metal-poor stars across 6 classical dwarf galaxies around the Milky Way, and 1604 stars across 19 globular clusters. We calibrate the reported abundance errors in bins of signal-to-noise and [Fe/H] using a high-fidelity halo dataset. Applying these calibrated errors to the APOGEE data, we find small amplitudes of average intrinsic abundance scatter in dwarf galaxies ranging from 0.03 -- 0.09 dex with a median value of 0.047 dex. For the globular clusters, we find intrinsic scatters ranging from 0.01 -- 0.11 dex, with particularly high scatter for Al and O. Our measurements of intrinsic abundance scatter place important upper bounds, which are limited by our calibration, on the intrinsic scatter in these systems, as well as constraints on their underlying star formation history and mixing, that we can look to simulations to interpret.
@@ 1 Introduction
Quenched dwarf galaxies that are bound to and inhabit our stellar halo serve as laboratories for studying snapshots of the early chemical evolution of the universe. With average stellar metallicities typically $\rm{[Fe/H]}<-1$ (Kirby et al., 2008, 2013; Geha et al., 2013; Li et al., 2018; Fu et al., 2023), their oldest, most metal-poor stars contain the chemical fingerprints of the very first generations of stars, including Population III (Frebel & Bromm, 2012; Ji et al., 2015; Hartwig et al., 2018, 2019). An understanding of the contribution to chemical enrichment from Population II stars is also necessary for placing constraints on the chemical evolution of the early universe. To decode the signatures of the chemical ancestry of dwarf galaxies, it is essential to understand the observational effects of the propagation of elements from various populations of stars and from a variety of nucleosynthetic sources throughout the interstellar medium (ISM; Ji et al., 2015).
@@
The chemical history of a stellar population is observed through measurements of element abundances, as well as their intrinsic abundance scatter – the true scatter of the data, accounting for measurement uncertainties – which necessitates careful consideration of uncertainties on abundances. Previous studies have uncovered overall element abundance scatter in numerous abundances, both within and between dwarf galaxies (Hill et al., 2019; Ji et al., 2020, 2022), but there remains a need to resolve the intrinsic scatters of chemical abundances above the level of observational errors; a difficult task for even small high-resolution spectroscopic studies. Mid-resolution spectroscopic surveys [e.g. APOGEE (Majewski et al., 2017), GALAH (De Silva et al., 2015), Gaia RVS (Gaia Collaboration et al., 2016, 2023) and LAMOST (Cui et al., 2012)] afford the community with copious amounts of data, and target numerous stellar systems. Nonetheless, measurements of the intrinsic abundance scatter are made even more challenging by medium-resolution data having larger measurement uncertainties than data from higher signal-to-noise (S/N), higher resolution data. Additionally, medium-resolution data may have potentially inaccurate estimates of measurement uncertainty in some parameter spaces. An important question to ask, however, is if we can use large mid-resolution surveys, to test the the intrinsic scatter of metal-poor systems that have been targeted and furthermore if we can validate or if necessary, calibrate, the uncertainty estimates so that these are accurate.
@@
From a theoretical perspective, the intrinsic scatter of element abundances in any system of stars is dependent on a number of properties. This includes the efficiency of gas mixing and metal diffusion in the ISM (Krumholz & Ting, 2018; Escala et al., 2018; Emerick et al., 2020a; Ji et al., 2023), metal-retention and recycling through inflows and outflows (Emerick et al., 2018, Mead et al. in prep), stochastic star formation (Welsh et al., 2021; Pan & Kravtsov, 2023; Griffith et al., 2023), mass and metallicity dependence of nucleosynthetic sources (Muley et al., 2021), and the assembly history of the system (Genina et al., 2019; Patel et al., 2022). The assessment of the scatter of individual elements is one indirect approach to access these environmental parameters.
@@
Because different nucleosynthetic channels contribute to the enrichment of different elements, intrinsic scatter is also a metric by which we can study the contribution of different nucleosynthetic sources to the chemical evolution of different stellar systems, as well as the mass and metallicity dependence of yields. For example, light elements such as C, N, and F are primarily produced by low and intermediate mass stars on the asymptotic giant branch (AGB) (Karakas, 2010; Kobayashi et al., 2011). Notably, some element abundances such as C, N, and O are sensitive to the evolutionary state of the star, in particular, mixing that brings the products of the CNO-cycle to the photosphere during the RGB phase. The $\alpha$-elements (e.g. O [^1] O is also produced through the CNO cycle, and so not always considered an $\alpha$-element, but this contribution is subdominant for low-metallicity stars., Mg, Si, Ca, Ti) are primarily produced by massive stars and ejected by core-collapse (Type II) supernovae (SNe; Timmes et al., 1995; Kobayashi et al., 2006). About half of Fe-peak elements (e.g. Cr, Mn, Fe, Ni, Co) are produced in Type Ia supernovae (Kobayashi & Nomoto, 2009). Putting these together, the well-known [$\alpha$/Fe]-[Fe/H] relationship, which shows the delay in enrichment in Fe from Type Ia SNe versus the short-timescale production of $\alpha$-elements by Type II SNe, can be used to characterize the star formation rates in galaxies (e.g. Tinsley, 1980; Matteucci & Greggio, 1986; Hayden et al., 2015).
@@
Other elements, such as the odd-Z elements (e.g. Al, P) and neutron-capture elements (s- and r-process elements) depend on the excess of neutrons available. Odd-Z elements depend on the metallicity of the progenitor star as their formation requires a surplus of neutrons made from 22Ne during He-burning (Kobayashi et al., 2011). Strong s-process elements (Sr to Pb) are produced in low-mass AGB stars (Busso et al., 1999; Herwig, 2005; Karakas & Lattanzio, 2014), whereas weak s-process elements (between Fe and Sr) are produced in massive stars near solar metallicity (Pignatari et al., 2010). The formation mechanism for r-process elements is debated but neutron star mergers are a likely scenario (Lattimer & Schramm, 1974; Rosswog et al., 1999; Goriely et al., 2011; Wanajo et al., 2014).
@@
Kobayashi et al. (2020) developed galactic chemical evolution models to place constraints on the origin of elements as a function of time and galactic environment. Models such as this, along with an understanding of gas mixing, the ratio of nucleosynthetic sources, and stochasticity of star formation can be used to predict the origin and amount of scatter in dwarf galaxy observations. High-resolution star-by-star models with individual stellar feedback (e.g. Brauer et al. in prep; Andersson et al., 2023) are likewise uniquely poised to explore these physical drivers in the context of intrinsic abundance scatter.
In the solar neighborhood, Griffith et al. (2023, hereafter G23) measured element abundances and intrinsic abundance scatters for metal-poor stars for 12 elements ranging from 0.04 for [Cr/Fe] to 0.16 for [Na/Fe] dex. Because of their low metallicity, these stars are likely field halo stars. They notably show that these scatters can be produced by Nsimilar-$\sim$50 core-collapse supernovae, which is a number too small to accurately represent an initial mass function (IMF).
@@ 2.1 Data Cuts
For all objects, we implement a metallicity cut of $-2.1\leq$ [Fe/H] $\leq-1$ to obtain the metal-poor stars whose scatter we are interested in. For all but the halo stars, we identify likely members of each system using the assigned APOGEE MEMBERFLAG. Halo stars are identified as metal-poor stars with distance $d\leq 2500\rm{pc}$. As in G23, these stars are from the solar neighborhood, but due to their low-metallicity, are likely halo stars. We remove stars in known substructure from this sample including Gaia-Sausage-Enceladus (GSE) using the same mask as Horta et al. (2023a), and globular clusters using the value-added catalog from Schiavon et al. (2024). We also include a stricter cut on GSE stars by eliminating high eccentricity (e $>$ 0.7) stars, but this had no effect on our results compared to a sample that omits this cut.
To ensure the quality of the data, we remove all stars which were flagged with warnings generated by ASPCAP for their stellar parameters (STAR_BAD) or abundance measurements (M_H_BAD; ELEMFLAG). To exclude very low-quality spectra, we only use stars with a $\rm{S/N}\geq 25$, with the final sample for dSphs having a median $\rm{S/N}=61.22$. Lastly, we only include stars with $4000\leq\rm{T_{eff}}\leq 5000K$ (median $\rm{T_{eff}}=4434\,\rm{K}$ for dSphs) and $0.0\leq\rm{log}g\leq 3.0$ (median $\rm{log}g=0.99$ for dSphs), keeping the range wide enough to contain as much of the data as possible but narrow enough to avoid systematic variations of these quantities with [X/Fe] (e.g. Griffith et al., 2021; Weinberg et al., 2022). Our final sample includes 333 stars in dSphs, 1604 stars in GCs and 291 stars in the halo. The Appendix includes a summary of the resulting number of stars in our sample for each object we study. Because Bootes I has so few metal-poor stars at the required S/N, we exclude Bootes I from analysis.
@@ 2.2 Data Summary
We use chemical abundances derived by ASPCAP for the stars in our sample (García Pérez et al., 2016). We focus on the following elements: Al (light elements); O, Mg, Si, Ti ($\alpha$-elements); and Ni and Mn (Fe-peak elements). Figure 1 shows the abundance [X/Fe] and metallicity [Fe/H] measurements for stars in our dSph sample. Data points in Figure 1 are colored by $\rm{T_{eff}}$. Generally, we see negative trends in [X/Fe] with [Fe/H]. As we are interested in processes that influence scatter in abundances for stars at the same metallicity, or in other words, the scatter around the trend line, we fit and subtract a second order model to the data in the [Fe/H]-[X/Fe] plane. We also account for trends in $\rm{T_{eff}}$ that may be expected for elements such as O, which are sensitive to the evolutionary state of the star, by including a linear term in our model as we want the measured scatter to be independent of stellar parameters. We perform a multivariate linear regression over the full model (Equation 1) to fit coefficients $a-e$.
$\rm{\langle[X/Fe]\rangle}=\rm{a[Fe/H]}+\rm{b[Fe/H]}^{2}+\rm{cT_{eff}}+\rm{d([Fe/H]\times T_{eff})}+\rm{e}$ (1)
We find that the absolute gradients of Teff for this regression are on the order $10^{-3}$ dex K-1, ranging from $8\times 10^{-4}$ to $3.5\times 10^{-3}$. Although the gradients appear small, over a 500K range, which is typical for a given [X/H]-dSph pair in our sample, they suggest a large variation of $\sim$0.5 dex. However, this apparent variation may not represent the true gradient due to the influence of the cross term. To demonstrate that the effect of Teff is in fact small, we report that the average $\Delta$[X/Fe] over the range of Teff for each [X/Fe]-dSph pair is only 0.18 dex for a model that does not include Teff terms.
Returning to the model in Equation 1, the gradients for the cross term are a factor of $\sim 2$ less than those for Teff. We nonetheless include the term as it appears to remove correlations between [Fe/H] and Teff, which may be an artifact of the data processing steps taken to produce the abundance catalog or the lack of data in the low-metallicity regime. In Figure 2, we show the resulting residuals for the dSphs, revealing the scatter of our sample with [Fe/H]. We similarly do this for the Milky Way stellar halo and GC stars in our sample. This trend-subtracted data is used in all further analysis.
Figure 3: Example line windows for star 2M02383095-3456544 in Fornax, which has an S/N of 58.6. The combined spectrum is shown in black with the ASPCAP model in red. The line window is indicated by the green shading.
@@ 2.3 Spectra Quality Check
Metal-poor stars have weak metal lines which make abundance measurements uncertain. Therefore, we visually examine the spectra of the 333 stars from dwarf galaxies in our sample to assess the quality of the ASPCAP fit for spectral lines of elements we are using. We use the strongest spectral lines from the line list (Smith et al., 2021) used in the APOGEE abundance pipeline, for a total of 18 lines, with at least 2 lines inspected per element. Ultimately, we find that Al, O, Mg, Si, Ti [^2] Ti is labeled as ‘deviant’ from literature expectations in DR17, however visual inspection shows the lines are well-fit for the stars in our sample., and Ni show absorption features, which are generally well-fit by the model. We conclude that it is reasonable to proceed with our analysis using these abundance measurements. We also include Mn in our analysis, but note that these lines were not as well-fit, and that these abundance measurements should be treated with caution and assumed to be upper limits. Figure 3 shows a typical example of ASPCAP line fits for the stars and elements included in our sample. In addition to a visual inspection, we implement a quality cut to all abundance measurements whereby we calculate the $\chi^{2}$ for the ASPCAP best fit model for each line (i.e. $\chi^{2}=\Sigma\frac{(\rm{x-y})^{2}}{\rm{\delta x}^{2}}$) and remove individual abundance measurements that do not have at least one line with a reduced $\chi^{2}>6$ from consideration in the analysis.
@@ 3.1 APOGEE Data Intrinsic Scatter
APOGEE has errors on the level of intrinsic scatter reported in other literature, therefore we compare the average scatters calculated with the APOGEE data used in our work to average scatters of other data for dwarfs in the literature, namely, the Kirby et al. (2018, hereafter K18) dataset which carefully accounts for the random and systematic uncertainties. From this data set, we only consider dSphs with at least 10 observed stars. This dataset overlaps partially with the APOGEE dataset (Draco, Fornax, Sculptor, Sextans, and Ursa Minor) and includes 3 additional dSphs (Canes Venatici I, Leo I, and Leo II). Figure 4 also shows that the uncalibrated data for the dSphs in the APOGEE dataset have higher average intrinsic scatter than the average for those in K18. This difference is substantial, and ranges from 0.07 to 0.11 dex.
Similarly high intrinsic abundance scatters can be shown for field stars in the APOGEE halo data when compared to intrinsic abundance scatters calculated in the same way from G23 data. In Figure 5, we report the intrinsic abundance scatters in the halo for the same elements. These range from a minimum of 0.08 dex for [Mg/Fe] to a maximum of 0.15 dex for [O/Fe]. We also show that these values range from 0.01 up to 0.1 dex higher than intrinsic scatters for halo stars in G23. These differences call into question the accuracy of the APOGEE uncertainty estimates for the metal-poor stars.